After 10 years, Brian's account earned $600 in interest. If the interest rate (in decimal form) is 0.08, how much did Brian initially invested?
Answer by stanbon(75887)

After 10 years, Brian's account earned $600 in interest. If the interest rate (in decimal form) is 0.08, how much did Brian initially invested?
------
I = P*r*t
600 = P*0.08*10
----
60 = 0.08P
----
P = 60/0.08
-----
P = 6000/8
-----
P = $750.00
------------
Cheers,
Stan H.
-------------